# The FFT Strikes Back: An Efficient Alternative to Self-Attention

### v2

@@ 3.2 Method Description
$\mathbf{c}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i},$
and pass it through a multi-layer perceptron (MLP) to obtain a modulation tensor:
$\Delta\mathbf{W}=\operatorname{MLP}(\mathbf{c})\;\in\;\mathbb{R}^{n\times d}.$
We define the final filter as
$\mathbf{W}=\mathbf{W}_{\mathrm{base}}+\Delta\mathbf{W},$
where $\mathbf{W}_{\mathrm{base}}$ is a fixed base filter (often initialized to all ones). The adaptive filtering step is then
$\tilde{\mathbf{F}}=\mathbf{F}\odot\mathbf{W},$
which reweights the Fourier coefficients element-wise according to the global context.
@@ 3.6 Proof of Computational Complexity
  • Computing $\Delta\mathbf{W}$ from $\mathbf{c}$ via a small MLP adds $\mathcal{O}(1)$ per channel.
@@ 3.7.2 Energy Preservation via Parseval’s Theorem
3.7.3 Approximation of Self-Attention Mechanism
Self-attention can be seen as a weighted sum of tokens via
$\mathbf{A}_{ij}\;=\;\frac{\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})}{\sum_{j=1}^{n}\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})},$
leading to
$\mathbf{Y}_{\text{attn}}\;=\;\mathbf{A}\,\mathbf{X}.$
Our method replaces explicit pairwise interactions with frequency-domain multiplication. By the convolution theorem, this frequency-domain multiplication is equivalent to a convolution in the token domain. Under mild conditions on $\mathbf{W}$, this global, data-dependent convolution can approximate self-attention at significantly lower computational cost.
@@ 3.7.4 Role of Nonlinear Activation in Enhancing Expressivity
3.8 Proof of Expressivity: FFT as an Approximation of Self-Attention
Theorem 1.
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ be an input sequence, and let the self-attention mechanism be
$\mathbf{Y}_{\text{attn}}=\mathbf{A}\mathbf{X},\quad\text{where}\quad\mathbf{A}_{ij}=\frac{\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})}{\sum_{j=1}^{n}\exp(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})}.$
Consider the adaptive spectral filtering operation
$\mathbf{Y}=\operatorname{IFFT}\Bigl{(}\operatorname{modReLU}\bigl{(}\operatorname{FFT}(\mathbf{X})\odot\mathbf{W}\bigr{)}\Bigr{)},$
where $\mathbf{W}$ is a learnable frequency-domain filter and $\operatorname{modReLU}$ is a complex-valued nonlinear activation. Then, under mild regularity conditions on $\mathbf{W}$ and the activation, there exists a parameterization such that
$\mathbf{Y}\approx\mathbf{Y}_{\text{attn}},$
and the presence of the nonlinear activation extends the expressive capacity beyond that of purely linear self-attention.
Proof.
We outline the argument in several steps:
1) Unitary Transformation and Energy Preservation. The DFT matrix $\mathbf{F}_{n}$ satisfies the unitary property (up to scaling),
$\mathbf{F}_{n}^{*}\mathbf{F}_{n}=n\mathbf{I},$
implying Parseval’s theorem, which preserves the norm of $\mathbf{X}$ across the frequency transform.
2) Frequency-Domain Filtering and Convolution Equivalence. The element-wise multiplication
$\operatorname{FFT}(\mathbf{X})\odot\mathbf{W}$
corresponds to a convolution in the token domain due to the convolution theorem. Defining $\mathbf{w}=\operatorname{IFFT}(\mathbf{W})$, we get
$\operatorname{IFFT}\Bigl{(}\operatorname{FFT}(\mathbf{X})\odot\mathbf{W}\Bigr{)}=\mathbf{X}\ast\mathbf{w}.$
3) Approximating Self-Attention via Convolution Kernels. Self-attention can be interpreted as a learnable, global aggregation function. Convolution kernels, particularly those conditioned on input features, can approximate a wide variety of such functions. Hence, suitable choices of $\mathbf{W}$ allow $\mathbf{X}\ast\mathbf{w}$ to approximate $\mathbf{Y}_{\text{attn}}$.
4) Nonlinear Activation for Enhanced Expressivity. Because data often require nonlinear modeling, we use $\operatorname{modReLU}$ on the complex coefficients:
$z=re^{i\theta}\quad\mapsto\quad\operatorname{modReLU}(z)=\begin{cases}(r+b)e^{i\theta},&\text{if }r+b>0,\\ 0,&\text{otherwise}.\end{cases}$
This nonlinearity, when applied element-wise in the frequency domain, enriches the effective convolution kernel beyond what a purely linear approach can achieve.
5) Conclusion. By choosing $\mathbf{W}$ and the bias $b$ in $\operatorname{modReLU}$ appropriately, we can approximate the self-attention operation with an FFT-based, $\mathcal{O}(n\log n)$ method while also harnessing the additional expressive potential afforded by nonlinear activation. ∎
@@ 4.2 ImageNet Classification
Next, we evaluate our FFTNetViT variants on the ImageNet classification task, comparing them to standard ViT (self-attention). Table 2 presents the FLOPs, parameter counts, and both Top-1 and Top-5 accuracy for each variant (Base, Large, Huge). We omit latency from the table for clarity and show it in a separate figure. Notably, FFTNetViT often achieves lower FLOPs than ViT for comparable model sizes, while maintaining strong accuracy. In most experiments, FNet lags behind both methods on ImageNet, so we focus on comparing our approach to the stronger self-attention baseline.
@@
FLOPs Params Top-1 (%) Top-5 (%) FLOPs Params Top-1 (%) Top-5 (%)
Base 22.64 76.33 79.6 94.9 36.65 86.57 79.4 94.8
Large 79.92 267.89 82.1 96.2 127.18 304.33 81.8 96.0
Huge 166.14 539.96 83.2 96.8 261.39 632.20 82.9 96.6
Table 2: Comparison of FFTNetViT vs. Standard ViT on ImageNet, grouped by variant. FLOPs are in GFLOPs and Params in millions.
Figure 1: Latency comparison of FFTNetViT vs. Standard ViT for varying batch sizes on ImageNet. FFTNetViT scales fastear than standard self-attention.

### v3

@@ 2.2 Fourier-Based Mixing
Beyond language and vision tasks, Fourier-based methods have also proven effective in modeling physical processes governed by partial differential equations. Fourier Neural Operator (FNO) Li et al. (2021) demonstrates that neural architectures parameterized in the Fourier domain can learn solution operators for parametric PDEs, achieving resolution-invariant performance and fast inference. Although primarily designed for PDE modeling and simulation, FNO nonetheless highlights the broader potential of leveraging Fourier representations for efficient token mixing in deep learning applications.
@@ 2.5 Orthogonal Matrix Decomposition Methods
Recently, there has also been significant progress in state space modeling approaches for sequence tasks, exemplified by S4 and its extensions. More streamlined variants such as the Simplified State Space layer (S5) Smith et al. (2023) have demonstrated that parameterizing the state space system with a diagonal transition matrix can enable efficient parallel scans over the sequence dimension while preserving high performance. These developments underscore the ongoing push to find alternatives to self-attention that can handle long sequence lengths without sacrificing expressiveness.
@@ 2.6 Adaptive Spectral Filtering in Context
Notably, our approach also differentiates itself from recent work such as Guibas et al. (2022), which introduces Adaptive Fourier Neural Operators (AFNO) for token mixing. While AFNO employs a block-diagonal structure for channel mixing and adaptive weight sharing, our method takes a distinct approach. Specifically, we compute a global context vector that directly modulates the Fourier coefficients via learnable scaling and bias parameters, followed by a nonlinear modReLU activation. This design not only preserves the computational advantages of FFT-based mixing but also enhances expressivity by dynamically emphasizing salient frequency bands based on the input. As such, our adaptive spectral filtering strategy offers a complementary alternative to AFNO, merging efficiency with adaptive, context-dependent feature modulation.
@@ 3.2 Method Description
$\mathbf{c}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{X}_{i}.$
This context is fed into a multi-layer perceptron that outputs two sets of modulation parameters defined over the frequency bins and attention heads: a multiplicative adaptive scale $\Delta\mathbf{s}$ and an additive adaptive bias $\Delta\mathbf{b}$, with $\Delta\mathbf{s},\,\Delta\mathbf{b}\in\mathbb{R}^{H\times f\times 1}$ (where $H$ is the number of heads and $f=\lfloor\tfrac{n}{2}\rfloor+1$ denotes the frequency bins). With a fixed base filter $\mathbf{W}_{\mathrm{base}}$ and a base bias $\mathbf{b}_{\mathrm{base}}$, the effective parameters are defined as:
$\mathbf{W}=\mathbf{W}_{\mathrm{base}}\odot\bigl{(}1+\Delta\mathbf{s}\bigr{)},\qquad\mathbf{b}=\mathbf{b}_{\mathrm{base}}+\Delta\mathbf{b}.$
The adaptive filtering step is then performed in the frequency domain:
$\tilde{\mathbf{F}}=\mathbf{F}\odot\mathbf{W}+\mathbf{b},$
which reweights and shifts the Fourier coefficients according to the global context.
@@ 3.6 Proof of Computational Complexity
  • Computing the adaptive modulation parameters via the MLP introduces negligible overhead.
@@ 3.9.1 Computational Efficiency
3.9.2 Mitigation Strategies for Adaptive MLP Complexity
The adaptive MLP in our spectral attention module generates modulation parameters for each frequency bin, where the number of bins is approximately $\frac{n}{2}+1$. Consequently, its output size scales linearly with the sequence length $n$, which may become a computational bottleneck for long sequences. We discuss several strategies to mitigate this issue:
  • Downsampling Frequency Bins: One effective approach is to reduce the number of frequency bins by applying pooling or subsampling techniques to the frequency-domain representation before feeding it into the adaptive MLP. This reduction preserves the essential spectral information while lowering the computational cost.
  • Parameter Sharing: Instead of generating independent modulation parameters for each frequency bin, the model can share parameters across groups of adjacent bins. This leverages the redundancy among nearby frequency components and significantly reduces the total number of parameters.
  • Low-Rank Approximations: By projecting the high-dimensional frequency representation into a lower-dimensional subspace using low-rank factorization, the adaptive MLP’s output size can be decreased. This strategy maintains the critical spectral features while reducing both the parameter count and computational overhead.
These techniques provide practical avenues to balance the trade-off between the expressive power of the adaptive MLP and its computational efficiency, especially when dealing with very long sequences.
@@ 4 Experiments
4.1 Training Scheme
We train all models using a combination of standard and advanced data augmentations (e.g., random crops, color jitter, random erasing, etc.) alongside mixup and cutmix strategies for better generalization. The optimization leverages AdamW with a cosine learning-rate schedule and a brief warmup phase. We also apply gradient clipping and optionally employ an exponential moving average (EMA) of parameters. Full details of these hyperparameters and augmentations are available in the accompanying training script.
@@ 4.3 ImageNet Classification
Next, we evaluate our FFTNetViT variants on the ImageNet classification task, comparing them to standard ViT (self-attention). Table 2 presents the FLOPs and both Top-1 and Top-5 accuracy for each variant (Base, Large, Huge). We omit latency from the table for clarity and show it in a separate figure. Notably, FFTNetViT often achieves lower FLOPs than ViT for comparable model sizes, while maintaining strong accuracy. In most experiments, FNet lags behind both methods on ImageNet, so we focus on comparing our approach to the stronger self-attention baseline.
@@
FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%)
Base 22.64 79.6 94.9 36.65 79.4 94.8
Large 79.92 82.1 96.2 127.18 81.8 96.0
Huge 166.14 83.2 96.8 261.39 82.9 96.6
Table 2: Comparison of FFTNetViT vs. Standard ViT on ImageNet, grouped by variant. FLOPs are in GFLOPs.
Figure 1: Latency comparison of FFTNetViT vs. Standard ViT for varying batch sizes on ImageNet. FFTNetViT scales faster than standard self-attention.